The pH of a 0.6M solution of a weak acid, HA (aq), is 1.98 at 25oC. What is the value of K?
a. 6.53 b. 1.86 x 10-4 c. 1.75 x 10-3 d. 1.8 x 10-2
1
Expert's answer
2020-05-05T03:40:25-0400

HA <-> H+ + A-

If x of HA have dissociated then

K = "{\frac {[H^+][A^-]}{[HA]}}={\frac {x*x}{0.6-x}}"

[H+] = [A-] = x = 10-pH = 10-1.98 = 0.0105

"{\frac {[H^+][A^-]}{[HA]}}={\frac {x*x}{0.6-x}}={\frac {0.0105*0.0105}{0.6-0.0105}}={\frac {0.00011}{0.5895}}=1.866*10^{-4}\approx1.86*10^{-4}"

Answer b 1.86 x 10-4